A shipment of 21 computer keyboards and 33 monitors cost $35,564.25. Another
shipment of 41 keyboards and 36 monitors cost $49,172.50. find the cost of each
keyboard and each monitor.

Answer by TimothyLamb(4379)

A = 35564.25
B = 49172.50
21k + 33m = A ... eqn 1
41k + 36m = B ... eqn 2
k = (B - 36m)/41
---
21( (B - 36m)/41 ) + 33m = A ... eqn 3
Answer:
you solve equation 3 for m then plug m into equation 1 and solve for k
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php